# Remasking Discrete Diffusion Models with Inference-Time Scaling

### v1

@@ 4.2 Design Strategies for ‘Turning On’ ReMDM
Loop  In this strategy, we set both $t_{on},t_{off}\in(0,1]$. Furthermore, in the range when ReMDM is activated, we modify the noise schedule to be constant, such that $\alpha_{t}=\alpha(t_{on}).$ As shown in Figure 2, this divides the sampling process into three phases. In the first phase, the model generates tokens without remasking ($\sigma_{t}=0$, i.e., using MDLM). In the second phase, we hold $\alpha$ constant (i.e., $\alpha_{s}=\alpha_{t}$), and the model can ‘correct potential mistakes’ by remasking and predicting a fixed proportion of the generated tokens in a loop. Finally, in the third phase, we let the model predict any remaining unmasked tokens using the MDLM posterior. We denote this strategy as “ReMDM-loop.” Note that we can also change the amount of computation spent in each of the three phases by taking non-fixed-width steps.
Figure 2: Illustration of $\alpha_{t}$ and $\sigma_{t}$ schedules in the ReMDM-loop.
@@ 4.3 Comparison with Discrete Predictor-Corrector Samplers
Previous works, e.g., the forward-backward (FB; Campbell et al. (2024)) and discrete flow matching (DFM; Gat et al. (2024)) correctors, propose to tackle the failure to remask property with discrete predictor-corrector samplers, a special type of discrete diffusion samplers that decompose a single sampling step into one predictor step followed by a certain number of corrector steps that remediate possible mistakes without changing the marginals. Here, we demonstrate that these methods are either a special case or a reparameterization of ReMDM sampling by first noting the following result (see Appendix A.5 for the proof):
@@
Since the only constraint on ReMDM’s sampler is the bound on $\sigma_{t}$, ReMDM offers a general framework for the design of remasking samplers for discrete diffusion models. We now formalize the generality of ReMDM  (proofs provided in Appendices A.6 & A.7).
@@
The FB corrector (Campbell et al., 2022) on MDLM is a special case of ReMDM where $\sigma_{t}=\frac{\alpha_{s}-\alpha_{t}}{\alpha_{t}}$.
@@
The DFM corrector (Gat et al., 2024) on MDLM is a reparameterization of ReMDM where $\sigma_{t}=\frac{\beta_{t}(\alpha_{s}-\alpha_{t})}{\alpha_{t}}$. $\beta_{t}\in\mathbb{R}$ denotes the corrector schecule.
@@ 5.1.1 Text Generation
Experimental Setup  We test the text generation capability of ReMDM with unconditional generation from models trained on OpenWebText (OWT; Gokaslan & Cohen (2019)). The OWT dataset was tokenized using the gpt-2 tokenizer (Radford et al., 2019) and sequences were wrapped to a max length of $L=1024$. Our baselines consist of pretrained AR, SEDD (Lou et al., 2024), and MDLM models with checkpoints taken from the implementation in Sahoo et al. (2024). We also compare to the MDLM pre-trained models with the FB (Campbell et al., 2022) and DFM (Gat et al., 2024) correctors during inference. For our method, we also reuse the MDLM checkpoint and experiment with different ReMDM remasking schedules, specifically the max-capped, rescaled, confidence-based, and loop schedules (see Appendix E.2.3 for the full set of results). We generate 5,000 samples from each model / sampler and evaluate methods using the MAUVE score (Liu et al., 2021; Pillutla et al., 2021), as this metric balances sample quality and diversity. We also report the perplexity of the generated sequences under a pre-trained GPT-2 Large model (Gen PPL.) to measure quality and average sequence entropy to reflect diversity of generated sequences as in (Zheng et al., 2024).
@@
Floating-Point Precision  As indicated in Zheng et al. (2024), previous masked diffusion models (Lou et al., 2024; Sahoo et al., 2024) report sampling with 32-bit floating point precision, which was found to significantly curb the diversity of generated samples. We therefore use 64-bit floating point for all text-sampling experiments.
@@
Nucleus Sampling  We observe that nucleus sampling, (Holtzman et al., 2019) is critical for generating high-quality text sequences. We therefore use this strategy (with top-p = 0.9) for all models (except SEDD—it doesn’t output logits).
@@
Results  In Table 1, we report results for the best ReMDM setting. For inference time scaling results ($T\geq 1024$), we use the max-capped schedule ($\eta_{cap}=0.02)$ in conjunction with the loop strategy ($t_{on}=0.55,t_{off}=0.05$, and $\alpha(t_{on})=0.9$ held constant in the ReMDM-loop). For faster sampling ($T<1024$), we use the max-capped schedule ($\eta_{cap}=0.04$) on its own.
@@
As shown in Table 1, ReMDM scales favorably with inference time compute, achieving a 15.62$\times$ MAUVE score compared to the masked diffusion models and a 2.23$\times$ MAUVE score compared to MDLM with corrector samplers. In contrast, masked diffusion models scale poorly with $T$ and corrector sampler methods saturate when $T$ is large.
We also explore the scenario of faster sampling (i.e., $T<L=1024$). Even with more limited computational budget, ReMDM is able to generate sequences with higher quality than the baselines.
@@
Ablation: Function of Each Component  In Figure 3, we sequentially ablate each building block of ReMDM and report the MAUVE score of samples. Starting with MDLM, we see that each of our proposed sampling improvements increases the MAUVE score, with the largest improvement coming from remasking ability of ReMDM.
Figure 3: Effect of ReMDM components on OWT generation quality. Inference-time scaling with $T\in\{1024,2048,4096\}$.
Ablation: Importance of Nucleus Sampling  In Appendix E.2.2, we present the generation quality of models without nucleus sampling. We find that discrete diffusion models suffer from poor generation quality, specifically due to failure to effectively reduce generative perplexity when nucleus sampling is removed. These results demonstrate the importance of nucleus sampling for discrete diffusion models in text generation.
Ablation: Sensitivity to $\eta$  In Appendix E.2.4, we show the parameter search results of ReMDM-cap, ReMDM-loop, and ReMDM-rescale. For larger $T$, as $\eta$ increases, the generative quality tends to first increase and then decrease.
@@
Table 1: ReMDM improves sample quality in the case of inference-time scaling and faster sampling. ReMDM outperforms state-of-the-art masked diffusion models (SEDD; Lou et al. (2024), MDLM; Sahoo et al. (2024)) and masked diffusion models with corrector samplers such as Forward-Backward (FB; Campbell et al. (2022)) and Discrete Flow Matching (DFM; Gat et al. (2024)) corrector samplers. indicates nucleus sampling. For each $T$, the best diffusion MAUVE score is bolded.
@@ 5.1.2 Image Generation
Experimental Setup  We use a pretrained MaskGiT model (Chang et al., 2022) that was trained on ImageNet (Deng et al., 2009) samples with $256\times 256$ pixels. The images were patchified and flattened to a sequence of $L=256$ and encoded according to a codebook with $1024$ tokens (Esser et al., 2021). See Chang et al. (2022) for more details about the model and training settings. We compare ReMDM’s sampler to the original MaskGiT sampling and to MDLM. Note that MaskGiT does not follow an ancestral sampling method but rather relies on an effective heuristic where masked tokens are decoded proportional to model confidence. However, MaskGiT is still restricted by the failure to remask property. For each sampler, we conditionally generate 50,000 images, with class labels randomly chosen from the 1,000 ImageNet categories, and we measure sample quality using Fréchet Inception Distance (FID; Heusel et al. (2017)) and Inception Score (IS; Salimans et al. (2016)). For all models, we explore the effect of scaling inference compute and tuning sampling temperature.
Results  We report the best results for each method in Table 2 (see Appendix E.3 for the more complete hyperparameter search). For MaskGiT, we found best results using no temperature, and for MDLM and ReMDM, we use a temperature of $0.8$. For our sampler, we report the ReMDM-rescale strategy with $\eta_{rescale}=0.05.$ Although for the smallest $T$ decoding setting MaskGiT outperforms the other methods, we see that ReMDM has the best scaling, producing the highest quality images of any model at $T=64$.
Ablation: Max-capped vs. Rescaled schedules  In Appendix E.3, we present the results from the full ReMDM parameter search for this experiment. We find that here ReMDM-rescale slightly outperforms ReMDM-cap and that for both increasing their corresponding $\eta$ parameters within the tested values leads to improved results.
Table 2: ReMDM produces the highest quality images. Values reflect FID / IS for varying $T$ on discretized ImageNet conditional generation. For each metric and $T$, the best value is bolded.
Metric Sampler $T=16$ $T=32$ $T=64$
@@ 5.2 ReMDM Improves Guidance
Experimental Setup  We follow the setup from Schiff et al. (2024) to explore controlled small molecule generation. Specifically, we use the QM9 dataset (Ruddigkeit et al., 2012; Ramakrishnan et al., 2014) of $\sim$133k molecules and their character-based SMILES string representations (Weininger, 1988). Sequences were tokenized using a regular expression method (Schwaller et al., 2019) and padded to a maximum length of $L=32.$ We use the D-CFG and D-CBG methods defined in Schiff et al. (2024) to conditionally generate molecules with higher ring counts (greater ${}^{\text{th}}$ percentile in the original dataset). We vary the guidance strength $\gamma\in\{1,2,3,4,5\}$ and scale inference steps $T\in\{32,64,128\}$. Our baselines consist of an AR model guided using either CFG or the popular classifier-based FUDGE method (Yang & Klein, 2021), MDLM, and the uniform diffusion language model (UDLM) proposed in Schiff et al. (2024), which was also introduced to alleviate the no remasking limitation of absorbing state diffusion. For ReMDM sampling, we use pretrained MDLM models. We explore the various strategies defined in Section 4 (see Appendix E.4 for full search results). After generating 1,024 sequences, we use the RDKit library (Landrum et al., 2013) to determine whether sequences are valid and compute novelty of the generated sequences (number of unique valid sequences that do not appear in the original QM9 dataset) and mean ring count of novel sequences.
@@
Results  In Figure 4, we display the trade-off that comes from increasing guidance strength $\gamma$. We only visualize results for samples that had at least 50 novel sequences. For both forms of guidance, CFG and CBG, ReMDM outperforms AR and diffusion approaches, pushing the novelty-property maximization frontier beyond that of the baseline methods. Additionally, ReMDM scales favorably with more inference-time compute, seen by the curves for larger $T$ dominating those for smaller $T$. For ReMDM, we found the best strategy to be a combination of the ReMDM-rescale (with $\eta_{rescale}=0.9$) and ReMDM-conf schedules. For calculating confidence scores, we use the log-probabilities of the conditional denoising network. For D-CFG, we use the loop strategy, with $t_{on}=0.25,t_{off}=0.125$. In addition to these findings, in Appendix E.4.1, we present experimental results for maximizing a different property of interest, drug-likeness (QED; Bickerton et al. (2012)).
@@
Figure 4: ReMDM improves steer-ability by extending the novelty-property maximization frontier. Controlled generation for ring count maximization on QM9 dataset with varying inference compute $T$ and guidance strength $\gamma$. (Left) Discrete classifier-free guidance (D-CFG). (Right) Discrete classifier-based guidance (D-CBG) and FUDGE for AR.
Ablation: Sensitivity to $\eta_{rescale}$  In Appendix E.4, we report extensive parameter sweeps for ReMDM. We find that larger $\eta_{rescale}$ values tend to perform better than smaller ones and relatively little sensitivity in the range $\eta_{rescale}\geq 0.5$ (see Appendix E.4.2).
@@
Ablation: Importance of using Confidence-based Schedule  Our parameter tuning revealed that using ReMDM-rescale in conjunction with the confidence-based scheduler led to improved results (see Appendix E.4.3).
Ablation: Tuning Switch / Loop Parameters  We observe a trend where for D-CFG results improve when the switch / loop start parameter ($t_{switch}$ / $t_{on}$) are smaller, i.e., we turn on ReMDM later in the decoding process. In contrast, for D-CBG, when maximizing ring count, we observe the opposite trend, where, generally, starting ReMDM earlier in the sampling process improves results (see Appendix E.4.4).
6 Discussion and Related Works

### v2

@@ 4.2 Design Strategies for ‘Turning On’ ReMDM
Loop  In this strategy, we set both $t_{on},t_{off}\in(0,1]$. Furthermore, in the range when ReMDM is activated, we modify the noise schedule to be constant, such that $\alpha_{t}=\alpha(t_{on}).$ As shown in Figure 2, this divides the sampling process into three phases. In the first phase, the model generates tokens without remasking ($\sigma_{t}=0$, i.e., using MDLM). In the second phase, we hold $\alpha$ constant (i.e., $\alpha_{s}=\alpha_{t}$), and the model is asked to remask and predict a proportion of the generated tokens in a loop. In particular, one can use any ReMDM strategy introduced in Section 4.1 in this phase. Intuitively, if a ‘bad’ token is remasked, it will likely be replaced with a ‘good’ token due to the abundant context. Even if a ‘good’ token happens to be remasked, since the signal-to-noise ratio is designed to be high in this portion of the generation process, it will likely be re-decoded as other (if not the same) ‘good’ tokens. In this way, ReMDM-loop corrects the ‘bad’ tokens and maintains the ‘good’ tokens, i.e., fixes mistakes. Finally, in the third phase, we let the model predict any remaining unmasked tokens using the MDLM posterior. We denote this strategy as “ReMDM-loop.”
Figure 2: ReMDM-loop illustration.
@@ 4.3 Comparison with Discrete Predictor-Corrector Samplers
Previous works, e.g., the forward-backward (FB; Campbell et al. [6]) and discrete flow matching (DFM; Gat et al. [13]) correctors, propose to tackle the failure to remask property with discrete predictor-corrector samplers, a special type of discrete diffusion samplers that decompose a single sampling step into one predictor step followed by a certain number of corrector steps that remediate possible mistakes without changing the marginals. Here, we demonstrate that these methods are special cases of ReMDM sampling by first noting the following result (see Appendix A.5 for the proof):
@@
We now formalize the generality of ReMDM  (proofs provided in Appendices A.6, A.7, and A.8).
@@
The FB corrector [5] on MDLM is a special case of ReMDM where $\sigma_{t}=\frac{\alpha_{s}-\alpha_{t}}{\alpha_{t}}$.
@@
A DFM corrector [13] on MDLM can be converted to a ReMDM sampler where $\sigma_{t}=\frac{\beta_{t}(\alpha_{s}-\alpha_{t})}{\alpha_{t}}$. $\beta_{t}\in\mathbb{R}$ denotes the DFM corrector schedule.
Proposition 4.4.
The ReMDM sampler is more general than DFM corrector, since only the former can accommodate noise schedules with constant $\alpha_{t}$ for some range $[t,t-\Delta t]$ where $\Delta t>0$.
@@ 5.1.1 Text Generation
Experimental Setup  We test the text generation capability of ReMDM with unconditional generation from models trained on OpenWebText (OWT; Gokaslan & Cohen [14]). The OWT dataset was tokenized using the gpt-2 tokenizer [35] and sequences were wrapped to a max length of $L=1024$. Our baselines include AR, SEDD [26], MDLM, and MDLM in conjunction with the FB [5] and DFM [13] correctors during inference. For all the experiments, we reuse the checkpoints trained by Sahoo et al. [38] for fair comparison. Following Pillutla [33], we generate 5,000 samples from each model/sampler and compute the MAUVE score [25, 34], as this metric balances sample quality and diversity.
Generative Perplexity vs. MAUVE It is common practice to report generative perplexity (Gen PPL.) as a quality metric for text generation. However, for corrector samplers such as ReMDM, we find this metric can be uninformative. In line with previous work [63], we observe that one can get arbitrarily low Gen PPL. by tuning corrector schedule hyperparameters (see Appendix E.2.2) to achieve low Gen PPL. while sacrificing sentence entropy/diversity, with an extreme case of generating a sequence of repetitive words like “love love love…" and getting a Gen PPL. of 1.5 (see Appendix F.1). Given this phenomenon, we choose MAUVE as the primary quality metric since MAUVE measures the balance between generative perplexity and diversity.
@@
Floating-Point Precision  As indicated in Zheng et al. [63], previous masked diffusion models [26, 38] report sampling with 32-bit floating point precision, which was found to significantly curb the diversity of generated samples. We therefore use 64-bit floating point for all text-sampling experiments.
@@
Nucleus Sampling  We observe that nucleus sampling [19] is critical to generating high-quality text sequences. We therefore use this strategy (with top-p = 0.9) for all models (except SEDD, as it does not output logits).
@@
Results  In Table 1, we report results for inference-time scaling ($T\geq L$) and faster sampling ($T<L$). Note that we report Gen PPL. and entropy for completeness and that lower Gen PPL. does not always correlate with better quality, as noted above. For inference-time scaling, we use the max-capped schedule ($\eta_{cap}=0.02)$ in conjunction with the loop strategy ($t_{on}=0.55,t_{off}=0.05$, and $\alpha(t_{on})=0.9$). For faster sampling, we use the max-capped schedule ($\eta_{cap}=0.04$) on its own.
@@
As shown in Table 1, ReMDM scales favorably with inference-time compute, achieving a 15.62$\times$ MAUVE score compared to masked diffusion models and a 2.23$\times$ MAUVE score compared to MDLM with corrector samplers. In contrast, masked diffusion models scale poorly with $T$, and corrector sampler methods saturate when $T$ is large. For faster sampling where a more limited computational budget is available, ReMDM is able to maintain sample quality better than baselines.
@@
Ablation: Function of Each Component  In Figure 3, we sequentially ablate each building block of ReMDM. Starting with MDLM, we see that each of our proposed sampling improvements increases the MAUVE score, with the largest improvement coming from the remasking ability of ReMDM.
@@
Table 1: ReMDM improves sample quality in inference-time scaling and faster sampling on OWT. ReMDM outperforms state-of-the-art masked diffusion models (SEDD; Lou et al. [26], MDLM; Sahoo et al. [38]) and masked diffusion models with corrector samplers such as Forward-Backward (FB; Campbell et al. [5]) and Discrete Flow Matching (DFM; Gat et al. [13]) corrector samplers. indicates nucleus sampling. For each $T$, the best diffusion MAUVE score is bolded.
@@ 5.1.2 Image Generation
Experimental Setup  We test ReMDM’s class-conditioned image generation ability. Concretely, we use a pretrained MaskGiT [7] that was trained on ImageNet [8] samples with $256\times 256$ pixels. The images were patchified and flattened to a sequence of $L=256$ and encoded according to a codebook with $1024$ tokens [10]. See Chang et al. [7] for more details about the model and training settings. We compare ReMDM’s sampler to the original MaskGiT sampling and to MDLM. Note that MaskGiT does not follow an ancestral sampling method but rather relies on an effective heuristic where masked tokens are decoded proportional to model confidence. However, MaskGiT is still restricted by the failure to remask property. For each sampler, we conditionally generate 50,000 images, with class labels randomly chosen from the 1,000 ImageNet categories, and we measure sample quality using Fréchet Inception Distance (FID; Heusel et al. [16]) and Inception Score (IS; Salimans et al. [42]).
Table 2: ReMDM produces the highest quality images on ImageNet conditional generation. For each metric and $T$, the best value is bolded.
Sampler $T=16$ $T=32$ $T=64$
@@
Results  We report the best results for each method in Table 2 (temperature=$1.0$ for MaskGiT, and temperature=$0.8$ for MDLM and ReMDM). We see that ReMDM has the best scaling, producing the highest quality images of the three methods at $T=64$.
Ablation: Max-capped vs. Rescaled schedules  In Appendix E.3, we present the full ReMDM parameter search and find that ReMDM-rescale slightly outperforms ReMDM-cap and that for both increasing their corresponding $\eta$ parameters leads to improved results.
@@ 5.2 ReMDM Improves Guidance
Experimental Setup  We follow the setup from Schiff et al. [43] to explore controlled small molecule generation. Specifically, we use the QM9 dataset [37, 36] of $\sim$133k molecules and their character-based SMILES string representations [55]. Sequences were tokenized using a regular expression method [44] and padded to a maximum length of $L=32.$ We use the D-CFG and D-CBG methods defined in Schiff et al. [43] to conditionally generate molecules with higher ring counts (greater ${}^{\text{th}}$ percentile in the original dataset). We vary the guidance strength $\gamma\in\{1,2,3,4,5\}$ and scale inference steps $T\in\{32,64,128\}$. Our baselines consist of an AR model guided using either CFG or the popular classifier-based FUDGE method [58], MDLM, and the uniform diffusion language model (UDLM) proposed in Schiff et al. [43], which was also introduced to alleviate the no remasking limitation of masked diffusion. For ReMDM sampling, we use pretrained MDLM models. We explore the various strategies defined in Section 4 (see Appendix E.4 for full search results). After generating 1,024 sequences, we use the RDKit library [21] to determine whether sequences are valid and compute novelty of the generated sequences (number of unique valid sequences that do not appear in the original QM9 dataset) and mean ring count of novel sequences.
@@
Results  In Figure 4, we display the trade-off that comes from increasing guidance strength $\gamma$. We only visualize results for samples that had at least 50 novel sequences. For both forms of guidance, CFG and CBG, ReMDM outperforms AR and diffusion approaches, pushing the novelty-property maximization frontier beyond that of the baseline methods. Additionally, ReMDM scales favorably with more inference-time compute, seen by the curves for larger $T$ dominating those for smaller $T$. For ReMDM, we found the best strategy to be a combination of the ReMDM-rescale (with $\eta_{rescale}=0.9$) and ReMDM-conf schedules. For D-CFG, we use the loop strategy, with $t_{on}=0.25,t_{off}=0.125$. In addition to these findings, in Appendix E.4.1, we present experimental results for maximizing a different property of interest, drug-likeness (QED; Bickerton et al. [3]).
@@
Figure 4: ReMDM improves steer-ability by extending the novelty-property maximization frontier. Controlled generation for ring count maximization on QM9 dataset, varying compute $T$ and guidance strength $\gamma$. (Left) Discrete classifier-free guidance (D-CFG). (Right) Discrete classifier-based guidance (D-CBG) and FUDGE for AR.
@@
Ablation: Confidence-based Schedule  Ablations show that using ReMDM-rescale in conjunction with the confidence-based scheduler led to improved results (see Appendix E.4.3 for more details).
5.3 ReMDM Improves Benchmark Performance
Experimental Setup Recent diffusion large language models (dLLMs) such as LLaDA 8B [28] and Dream 7B [60] have demonstrated on par downstream task performance with autoregressive language models of the same parameter scale. However, these dLLMs still suffer from the failure to remask property. To test the effect of remasking sampling on dLLLMs, we take LLaDA 8B Instruct and apply DFM and ReMDM samplers (see Appendix D.4 for sampler details). Specifically, we use Countdown (bidirectional reasoning) [59] and TruthfulQA (factual knowledge grasp) [24]. In Countdown, the model is given four integers and asked to develop equations where first three integers produce the fourth. In TruthfulQA, the model is asked 817 questions on factual knowledge. We calculate the ROUGE [23] scores between the model and correct answers, as well as the model and plausibly wrong answers. The difference between the right and wrong ROUGE scores is reported. For both of the tasks, we use max-length $L=32$ and no semi-autoregressive sampling [28]. See Appendix D.4 & F.3 for more setting details.
Results  Since the metric gains are not as significant as in Table 1, we repeat experiments with multiple random seeds to report mean values and 95% confidence intervals. In Table 3, LLaDA with ReMDM consistently performs the best and exceeds the original LLaDA with statistical significance.
Table 3: LLaDA with ReMDM performs the best on downstream tasks, and consistently exceeds the original LLaDA with statistical significance. For each metric, the best value is bolded.
Countdown TruthfulQA
Metric pass@1% $\Delta$ ROUGE 1/2/L
LLaDA 45.2±${}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\text{0.2}}}$ 27.1±${}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\text{0.4}}}$ / 30.1±${}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\text{0.4}}}$ / 27.2±${}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\text{0.4}}}$
LLaDA-DFM 44.8±${}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\text{0.2}}}$ 28.2±${}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\text{0.4}}}$ / 31.1±${}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\text{0.4}}}$ / 28.3±${}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\text{0.4}}}$
LLaDA-ReMDM 46.1±0.2 29.5±0.4 / 31.8±0.4 / 29.5±0.3
6 Discussion and Related Work